OpenStudy (anonymous):

INTEGRAL CALCULUS -Integral of logarithmic functions "Evaluate this integral of (x+3)dx/x+2

OpenStudy (anonymous):

I think you'll have to use long division first since both sides have the same power. The numerator's power must be < the denominator's power. :)

OpenStudy (anonymous):

can u answer it? i want to compare my answer.. please?

OpenStudy (anonymous):

alright, did you do the long division part?

OpenStudy (anonymous):

I got this after diving x+3 / x+2 : 1 + 1/x+2

OpenStudy (anonymous):

dividing*

OpenStudy (anonymous):

is this what you've got?

OpenStudy (anonymous):

my answer x+ln(x+2)+c is this correct?

OpenStudy (anonymous):

hold on , let me finish :)

OpenStudy (anonymous):

yes, excellent :)

OpenStudy (anonymous):

here another example can u answer this

OpenStudy (anonymous):

sure, I'll give it a try ^_^

OpenStudy (anonymous):

integral of xdx/x^2+4

OpenStudy (anonymous):

you can expand the denominator and you'll get the following:\[x/(x^2+4) = x/(x+2)^2\] then you'll have to do the partial fractions part which is : \[x/(x+2)^2 = A/(x+2) + B/(x+2)^2\]....etc correct me if I'm wrong

OpenStudy (anonymous):

you know how to integrate partial fractions, right? :)

OpenStudy (anonymous):

wers the answer?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

lol, I have hinted the way for you for you to proceed ^_^

OpenStudy (anonymous):

yes you know :) haven't you taken partial fractions?

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

ther's a much simpler way to do it :) hold on no need for partial fractions here

OpenStudy (anonymous):

ur answer? i want to compare my answer

OpenStudy (anonymous):

\[\int\limits_{} x/(x^2+4) dx\] let : \[u = x^2 +4\] \[du = 2x dx\] so : \[= 1/2 \int\limits_{} 1/u du\] \[= 1/2(\ln |u|) + c\] \[=[ \ln|x^2+4|] /2 + c\] sorry abt the complicated mess earlier ^_^

OpenStudy (anonymous):

is this what you've got?

OpenStudy (anonymous):

1/2ln(X^2+4)+C

OpenStudy (anonymous):

yes, well done ^_^

OpenStudy (anonymous):

i wish you can transfer your knowledge in calculus to me

OpenStudy (anonymous):

lol, I was so ignorant in it, but then worked hard to understand it. I took this in the beginning of the semester :), just be patient and practice more then you'll master it ^_^

OpenStudy (anonymous):

do you have facebook i wish we can talk together to practice my enlish

OpenStudy (anonymous):

english

OpenStudy (anonymous):

Unfortunately, I don't have FB, myspace or any of these stuff since I don't have time for them :( , I'm sorry.

OpenStudy (anonymous):

To perfect your english, read more, listen more and communicate more :) that's it

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

np :)